# Graph-Structured Topic Modeling for Documents with Spatial or Covariate Dependencies

### v1

@@ Appendix A Analysis of the initialization step
In this section, we show that the initialization in Algorithm 3 provides reasonable estimators of $U$ and $V$.
@@
Proof  We have, by definition of the initialization procedure:
$\widehat{V}_{0}=U_{K}(X^{\top}X-\frac{n}{N}\widehat{D}_{0}).$
Let $D_{0}$ denote the diagonal matrix where each entry $(D_{0})_{jj}$ is defined as: $(D_{0})_{jj}=\frac{1}{n}\sum_{i=1}^{n}M_{ij}.$ Let $\widehat{D}_{0}$ denote its empirical counterpart, that is, the diagonal matrix defined as: $\widehat{D}_{0}=\text{diag}(\frac{1}{n}\{\sum_{i=1}^{n}X_{ij}\}_{j\in[p]})$, so that $\mathbb{E}[\widehat{D}_{0}]=D_{0}$.
We have:
@@
[…]
@@
Thus $\mathbb{E}[{X^{\top}X}-\frac{n}{N}\widehat{D}_{0}]=(1-\frac{1}{N})M^{\top}M.$ We further note that $(1-\frac{1}{N})M^{\top}M=V\tilde{\Lambda}^{2}V$ with $\tilde{\Lambda}=(1-\frac{1}{N})\Lambda,$ so the eigenvectors of the matrix ${X^{\top}X}-\frac{n}{N}\widehat{D}_{0}$ can be considered as estimators of those of the matrix $M^{\top}M.$
By the Davis-Kahan theorem (Giraud, 2021):
@@
$\begin{split}\|\sin\Theta(V,\widehat{V}^{0})\|_{F}&\leq\frac{14C}{\lambda_{K}(M)^{2}}K\sqrt{\frac{n\log(n)}{N}}\end{split}$
@@
$\begin{split}\|\sin\Theta(V,\widehat{V}^{0})\|_{F}&\leq\frac{14C}{c^{2}}K^{2}\sqrt{\frac{\log(n)}{nN}}\\ \end{split}$
Note that the condition of $N$ defined in (12) in Theorem 2 ensures that $\|\sin\Theta(V,\widehat{V}^{0})\|_{F}<\frac{1}{2}$.
Appendix B Analysis of iterative graph-aligned denoising
@@
    We show that, by a simple readaptation of the proof by Klopp et al. (2021), our estimator—which simply plugs in our singular vector estimates in the procedure of Klopp et al. (2021)—yields a better estimate of the mixture matrix $W$.
@@
    Finally, we show that our estimator of the topic matrix $A$ yields a better error.
B.1 Analysis of the graph-regularized SVD procedure
@@
$L_{t}=\max\{\|\sin(\Theta(V,\widehat{V}^{t})\|_{F},\|\sin(\Theta(U,\widehat{U}^{t})\|_{F}\}.$ (25)
Our proof is a proof by recursion. We explicit the dependency of $L_{t}$ on the error at the previous iteration $L_{t-1}$, and show that $\{L_{t}\}_{t=1,\cdots,t_{\max}}$ forms a geometric series. To this end, we begin by analyzing the error of the denoised matrix $\bar{U}^{t}$, of which we later take an SVD to extract $\widehat{U}^{t}$.
B.1.1 Effect of the graph-denoising step
@@
$\bar{U}^{t}\in\arg\min_{U\in\mathbb{R}^{n\times K}}\|{U-X\widehat{V}^{t-1}}\|_{F}^{2}+\rho\|{\Gamma U}\|_{2,1}\$ (26)
@@
Lemma 1 (Error bound of Graph-aligned Denoising)
@@
$\|\bar{U}^{t}-\tilde{U}\|_{F}\leq C\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s}\sqrt{\lambda_{\max}}(1+L_{t-1})\right)$ (29)
Proof  By the KKT conditions:
@@
$\begin{split}\langle\tilde{Y}-\bar{U}^{t},\bar{U}^{t}\rangle&=\frac{\rho}{2}\langle\Gamma\bar{U}^{t},D\Gamma\bar{U}^{t}\rangle=\frac{\rho}{2}\|\Gamma\bar{U}^{t}\|_{21}\\ \text{and}\quad\forall U\in\mathbb{R}^{n\times p},\qquad\langle\tilde{Y}-\bar{U}^{t},U\rangle&=\frac{\rho}{2}\langle\Gamma U,D\Gamma\bar{U}^{t}\rangle\leq\frac{\rho}{2}\|\Gamma{U}\|_{21}\\ \end{split}$
@@
We turn to the control of the error term $\langle\tilde{Z},\bar{U}^{t}-\tilde{U}\rangle$ Using the decomposition of $I_{n}=\Pi\oplus^{\perp}\Gamma^{\dagger}\Gamma$, we have:
@@ B.1.1.1 Bound on (A) in (30)
We know that $\Pi_{[C_{l}]}=\frac{1}{n_{\mathcal{C}_{l}}}\mathbf{1}_{\mathcal{C}_{l}}\mathbf{1}_{\mathcal{C}_{l}}\top.$ Using Lemma 14, with probability at least $1-o(n^{-1})$:
@@
$\begin{split}\|\Delta\|_{F}^{2}&\leq\|\Pi\tilde{Z}\|_{F}\|\Delta\|_{F}+\frac{3\rho}{4}\|\Gamma\Delta\|_{21}\\ &\leq\|\Pi\tilde{Z}\|_{F}\|\Delta\|_{F}+\frac{3\rho}{4}\sqrt{s}\|\Gamma\Delta\|_{F}\\ &\leq\|\Pi\tilde{Z}\|_{F}\|\Delta\|_{F}+\frac{3\rho}{4}\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\|\Delta\|_{F}\\ \end{split}$ (31)
@@
$\begin{split}\|\tilde{U}-\bar{U}^{t}\|_{F}&\leq C_{1}\sqrt{\frac{n_{\mathcal{C}}K\log(n)}{N}}+3C_{2}\rho(\Gamma)\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\sqrt{\frac{K\log(n)}{N}}(1+L_{t-1})\\ &\leq C\sqrt{\frac{K\log(n)}{N}}(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}(1+L_{t-1}))\\ \end{split}$
B.1.2 Proof of Theorem 2
Proof  Let $L_{t}$ be the error at each iteration $t$:
$L_{t}=\max\{\|\sin(\Theta(V,\widehat{V}^{t})\|_{F},\|\sin(\Theta(U,\widehat{U}^{t})\|_{F}\}.$ (32)
B.1.2.1 Bound on $\|\sin(\Theta(U,\widehat{U}^{t})\|_{F}$.
@@ B.1.2 Proof of Theorem 2
Noting that $\widehat{U}^{t}$ is the matrix corresponding to the top $K$ left singular vectors of the matrix $\tilde{U}^{t}=(\tilde{U}^{t}-MV)+MV=(\tilde{U}^{t}-M\widehat{V}^{t}+M\widehat{V}^{t}-MV)+MV,$ by Theorem 1 of Cai and Zhang (2018) (Lemma  6 in the appendix):
@@
Thus, in this setting:
@@
B.1.2.2 Bound on $\|\sin\Theta(V,\widehat{V}^{t})\|_{F}$
@@
$\widehat{V}^{t}=SVD_{K}(X^{\top}\widehat{U}^{t}).$
By Theorem 1 of Cai and Zhang (2018) (summarized for our use case in Lemma 6 in Appendix C):
@@
$\begin{split}\lambda_{\min}(V^{\top}X^{\top}\widehat{U}^{t})&=\lambda_{\min}(V^{\top}M^{\top}\widehat{U}^{t}+V^{\top}Z^{\top}\widehat{U}^{t})\\ &\geq\lambda_{\min}(\Lambda U^{\top}\widehat{U}^{t})-\lambda_{\max}(V^{\top}Z^{\top}\widehat{U}^{t})\qquad\text{ (Weyl's inequality)}\\ &=\lambda_{\min}(\Lambda)\underbrace{\lambda_{\min}(U^{\top}\widehat{U}^{t})}_{=\sqrt{1-L_{t}^{2}}}-\|V^{\top}Z^{\top}\widehat{U}^{t}\|_{F}\\ &\geq\lambda_{\min}(\Lambda)\sqrt{1-L_{t}^{2}}-\|V^{\top}Z^{\top}\widehat{U}^{t}\|_{F}\end{split}$
@@
$\begin{split}n&\geq\frac{100C^{2}}{c^{2}}{\frac{K^{2}n\log(n)}{N}}\\ \implies N&\geq\frac{100C^{2}}{nc^{2}}{{K^{2}\log(n)}}\end{split}$ (35)
Then we can show that,
@@ B.1.2.3 Behavior of $L_{t}$
$\begin{split}\|\sin\Theta(U,\widehat{U}^{t})\|_{F}&\leq\frac{5/2C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\qquad+\frac{200/49C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}L_{t-1}\\ \|\sin\Theta(V,\widehat{V}^{t})\|_{F}&\leq\frac{5/2C}{\lambda_{K}(M)}K\sqrt{\frac{\log(n)}{N}}+\frac{200/49C}{\lambda_{K}(M)}\sqrt{\frac{Kn\log(n)}{N}}L_{t-1}\end{split}$
@@
$\begin{split}u&=\frac{5/2C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ r&=\frac{200/49C}{\lambda_{K}(M)}K\sqrt{\frac{\log(n)}{N}}\left(\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\vee\sqrt{n}\right)\end{split}$
where $r\leq\frac{1}{2},$ In particular, we want to find $t_{\max}$ such that $r^{t_{\max}}L_{0}\leq\frac{u}{1-r}$. Using $r\leq\frac{1}{2}$ (as previously shown) and that $L_{0}\leq\frac{1}{2}$,
@@
Denote $m_{1}=\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}$ and $m_{n}=\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\vee\sqrt{n}$. Lastly, once $t_{\max}$ is chosen as (38), the bound on $L_{t_{\max}}$ in (37) becomes,
@@
$\begin{split}L_{t_{\max}}&\leq\frac{2u}{1-r}\\ &=\frac{10C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\leq\frac{10C}{c}K\sqrt{\frac{\log(n)}{nN}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\leq\frac{1}{c}\end{split}$ (39)
where $c>1$ (Assumption 4). This concludes the proof.
B.2 Analysis of the Estimation of $W$

### v2

@@ Appendix B Analysis of the initialization step
In this section, we show that the initialization step of GpLSI provides reasonable estimators of $U$ and $V$.
@@
Proof.
Let $D_{0}$ denote the diagonal matrix where each entry $(D_{0})_{jj}$ is defined as: $(D_{0})_{jj}=\frac{1}{n}\sum_{i=1}^{n}M_{ij}.$ Let $\widehat{D}_{0}$ denote its empirical counterpart, that is, the diagonal matrix defined as: $\widehat{D}_{0}=\text{diag}(\frac{1}{n}\{\sum_{i=1}^{n}X_{ij}\}_{j\in[p]})$, so that $\mathbb{E}[\widehat{D}_{0}]=D_{0}$. We have, by definition of the initialization procedure:
$\widehat{V}_{0}=U_{K}(X^{\top}X-\frac{n}{N}\widehat{D}_{0}),$
where the notation $U_{K}(A)$ denotes the first $K$ left singular vectors of the matrix $A$.
We write $X=M+Z$, where $Z$ denotes some multinomial noise. E We have:
@@
as $Z$ is a centered version of $X$ ($Z=X-M$). Since each row $X_{i\cdot}$ is distributed as a Multinomial$(1,M_{i\cdot})$:
@@
Thus $\mathbb{E}[{X^{\top}X}-\frac{n}{N}\widehat{D}_{0}]=(1-\frac{1}{N})M^{\top}M.$ We further note that $(1-\frac{1}{N})M^{\top}M=V\tilde{\Lambda}^{2}V$ with $\tilde{\Lambda}=\sqrt{1-\frac{1}{N}}\Lambda,$ so the eigenvectors of the matrix ${X^{\top}X}-\frac{n}{N}\widehat{D}_{0}$ can be considered as estimators of those of the matrix $M^{\top}M.$
By the Davis-Kahan theorem (Giraud 2021):
@@
$\begin{split}\|\sin\Theta(V,\widehat{V}^{0})\|_{F}&\leq\frac{4C}{\lambda_{K}(M)^{2}}K\sqrt{\frac{n\log(n)}{N}}\end{split}$
@@
$\begin{split}\|\sin\Theta(V,\widehat{V}^{0})\|_{F}&\leq\frac{4C}{c^{2}}K^{2}\sqrt{\frac{\log(n)}{nN}}\\ \end{split}$
The condition on $N$ assumed in Theorem 2 ensures that $\|\sin\Theta(V,\widehat{V}^{0})\|_{F}<\frac{1}{2}$. ∎
Appendix C Analysis of iterative graph-aligned denoising
@@
    We show that, by a simple readaptation of the proof by Klopp et al. (2021), our estimator—which simply plugs in our singular vector estimates in their procedure —yields a better estimate of the mixture matrix $W$.
@@
    Finally, we show that our estimator of the topic matrix $A$ yields better error.
C.1 Analysis of the graph-regularized SVD procedure
@@
$L_{t}=\max\{\|\sin\Theta(V,\widehat{V}^{t})\|_{F},\|\sin\Theta(U,\widehat{U}^{t})\|_{F}\}.$ (24)
Our proof operates by recursion. We explicit the dependency of $L_{t}$ on the error at the previous iteration $L_{t-1}$, and show that $\{L_{t}\}_{t=1,\cdots,t_{\max}}$ forms a geometric series. To this end, we begin by analyzing the error of the denoised matrix $\bar{U}^{t}$, of which we later take an SVD to extract $\widehat{U}^{t}$.
@@
$\bar{U}^{t}\in\arg\min_{U\in\mathbb{R}^{n\times K}}\|{U-X\widehat{V}^{t-1}}\|_{F}^{2}+\rho\|{\Gamma U}\|_{21}\$ (25)
@@
Lemma 1 (Error bound of Graph-aligned Denoising).
@@
$\|\bar{U}^{t}-\tilde{U}\|_{F}\leq C\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s}\sqrt{\lambda_{\max}(L)}(1+L_{t-1})\right)$ (28)
where $L$ denotes the graph Laplacian.
Proof.
By the KKT conditions, the solution $\bar{U}^{t}$ of (25) verifies :
@@
$\begin{split}\langle\tilde{Y}-\bar{U}^{t},\bar{U}^{t}\rangle&=\frac{\rho}{2}\langle\Gamma\bar{U}^{t},D\Gamma\bar{U}^{t}\rangle=\frac{\rho}{2}\|\Gamma\bar{U}^{t}\|_{21}\\ \text{and}\quad\forall U\in\mathbb{R}^{n\times K},\qquad\langle\tilde{Y}-\bar{U}^{t},U\rangle&=\frac{\rho}{2}\langle\Gamma U,D\Gamma\bar{U}^{t}\rangle\leq\frac{\rho}{2}\|\Gamma{U}\|_{21}\\ \end{split}$
@@
We turn to the control of the error term $\langle\tilde{Z},\bar{U}^{t}-\tilde{U}\rangle$. Using the decomposition of $\mathbb{R}^{n}$ induced by the projection $\Gamma^{\dagger}\Gamma$ as $I_{n}=\Pi\oplus^{\perp}\Gamma^{\dagger}\Gamma$, we have:
@@
By Lemma 16, with probability at least $1-o(n^{-1})$:
@@
$\begin{split}\|\Delta\|_{F}^{2}&\leq\|\Pi\tilde{Z}\|_{F}\|\Delta\|_{F}+\frac{3\rho}{4}\|\Gamma\Delta\|_{21}\\ &\leq\|\Pi\tilde{Z}\|_{F}\|\Delta\|_{F}+\frac{3\rho}{4}\sqrt{s}\|\Gamma\Delta\|_{F}\\ &\leq\|\Pi\tilde{Z}\|_{F}\|\Delta\|_{F}+\frac{3\rho}{4}\sqrt{s}{\lambda_{\max}(\Gamma)}\|\Delta\|_{F}\\ \end{split}$ (30)
@@
$\begin{split}\|\tilde{U}-\bar{U}^{t}\|_{F}&\leq C_{1}\sqrt{\frac{n_{\mathcal{C}}K\log(n)}{N}}+3C_{2}\rho(\Gamma)\sqrt{s}{\lambda_{\max}(\Gamma)}\sqrt{\frac{K\log(n)}{N}}(1+L_{t-1})\\ &\leq C\sqrt{\frac{K\log(n)}{N}}(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s}{\lambda_{\max}(\Gamma)}(1+L_{t-1}))\\ \end{split}$
The result follows by noting that the Laplacian of the graph $L$ is linked to $\Gamma$ by $L=\Gamma^{\top}\Gamma$. ∎
Proof of Theorem 2
Proof.
Recall $L_{t}$, the error at each iteration $t$:
$L_{t}=\max\{\|\sin\Theta(V,\widehat{V}^{t})\|_{F},\|\sin\Theta(U,\widehat{U}^{t})\|_{F}\}.$ (31)
Bound on $\|\sin(\Theta(U,\widehat{U}^{t})\|_{F}$.
@@
Noting that $\widehat{U}^{t}$ is the matrix corresponding to the top $K$ left singular vectors of the matrix $\tilde{U}^{t}=(\tilde{U}^{t}-MV)+MV=(\tilde{U}^{t}-M\widehat{V}^{t}+M\widehat{V}^{t}-MV)+MV,$ by Theorem 1 of Cai & Zhang (2018) (which we rewrote in Lemma  8 of this manuscript to make it self-contained):
@@
which is satisfied under the condition (12) of $N$ in Theorem 2. Thus, in this setting:
@@
Bound on $\|\sin\Theta(V,\widehat{V}^{t})\|_{F}$
@@
$\widehat{V}^{t}=U_{K}(X^{\top}\widehat{U}^{t}).$
By Theorem 1 of Cai & Zhang (2018) (summarized for our use case in Lemma 8):
@@
$\begin{split}\lambda_{\min}(V^{\top}X^{\top}\widehat{U}^{t})&=\lambda_{\min}(V^{\top}M^{\top}\widehat{U}^{t}+V^{\top}Z^{\top}\widehat{U}^{t})\\ &\geq\lambda_{\min}(\Lambda U^{\top}\widehat{U}^{t})-\lambda_{\max}(V^{\top}Z^{\top}\widehat{U}^{t})\qquad\text{ (Weyl's inequality)}\\ &=\lambda_{K}(M)\underbrace{\lambda_{\min}(U^{\top}\widehat{U}^{t})}_{=\sqrt{1-L_{t}^{2}}}-\|V^{\top}Z^{\top}\widehat{U}^{t}\|_{F}\\ &\geq\lambda_{K}(M)\sqrt{1-L_{t}^{2}}-\|V^{\top}Z^{\top}\widehat{U}^{t}\|_{F}\end{split}$
@@
$\begin{split}n&\geq\frac{100C^{2}}{c^{2}}{\frac{K^{2}n\log(n)}{N}}\\ \implies N&\geq\frac{100C^{2}}{c^{2}}{{K^{2}\log(n)}}\end{split}$ (34)
which is satisfied under the condition (12) of $N$ in Theorem 2. Then we can show that,
@@
$\begin{split}\|\sin\Theta(U,\widehat{U}^{t})\|_{F}&\leq\frac{5/2\eta_{n}}{\lambda_{K}(M)}+\frac{200/49\delta_{n}}{\lambda_{K}(M)}L_{t-1}\\ &\leq\frac{5/2C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\qquad+\frac{200/49C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}L_{t-1}\\ \|\sin\Theta(V,\widehat{V}^{t})\|_{F}&\leq\frac{5/2\tilde{\eta}_{n}}{\lambda_{K}(M)}+\frac{200/49\tilde{\delta}_{n}}{\lambda_{K}(M)}L_{t-1}\\ &\leq\frac{5/2C}{\lambda_{K}(M)}K\sqrt{\frac{\log(n)}{N}}+\frac{200/49C}{\lambda_{K}(M)}\sqrt{\frac{Kn\log(n)}{N}}L_{t-1}\end{split}$
@@
$\begin{split}u&=\frac{5/2C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ r&=\frac{200/49C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\vee\sqrt{n}\right)\end{split}$
where $r\leq\frac{1}{2},$ In particular, we want to find $t_{\max}$ such that $r^{t_{\max}}L_{0}$ becomes small enough to satisfy $r^{t_{\max}}L_{0}\leq\frac{u}{1-r}$. Using $r\leq\frac{1}{2}$ (as previously shown) and that $L_{0}\leq\frac{1}{2}$,
@@
Lastly, once $t_{\max}$ is chosen as (37), the bound on $L_{t_{\max}}$ in (36) becomes,
@@
$\begin{split}L_{t_{\max}}&\leq\frac{2u}{1-r}\leq 4u\\ &=\frac{10C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\\ &\leq\frac{10C}{c}K\sqrt{\frac{\log(n)}{nN}}\left(\sqrt{n_{\mathcal{C}}}+\rho(\Gamma)\sqrt{s\lambda_{\max}(\Gamma)}\right)\end{split}$ (38)
This concludes the proof.
C.2 Comparison with One-step Graph-Aligned denoising
We also propose a fast one-step graph-aligned denoising of the matrix $X$ that could be an alternative of the iterative graph-aligned SVD in Step 1 of Section 2.3 of the main manuscript. We denoise the frequency matrix $X$ by the following optimization problem,
A SVD on the denoised matrix $\widehat{M}$ yields estimates of the singular values $U$ and $V$. Through extensive experiments with synthetic data, we find that one-step graph-aligned denoising provides more accurate estimates than pLSI but still falls short compared to the iterative graph-aligned denoising (GpLSI). We provide a theoretical upper bound on its error as well as its comparison to the error of pLSI where there is no graph-aligned denoising.
Algorithm 3 One-step Graph-aligned denoising
1:Input: Observation $X$, incidence matrix $\Gamma$
2:Output: Denoised singular vectors $\widehat{U}$ and $\widehat{V}$.
We begin by analyzing the one-step graph-aligned denoising, as proposed in Algorithm 3. We begin by reminding the reader that, in our proposed setting, the observed word frequencies in each document are assumed to follow a “signal + noise” model, $X=M+Z$ where the true probability $M$ is assumed to admit the following SVD decomposition:
$M=\mathbb{E}[X]=U\Lambda V^{\top}.$
Theorem 5.
Let the conditions of Theorem 3 hold. Let $\widehat{U}$ and $\widehat{V}$ be given as estimators obtained from Algorithm 3. Then, there exists a constant $C>0$, such that with probability at least $1-o(n^{-1})$,
Proof.
Let $\widehat{M}$ be the solution of (39):
Let $\Delta=\widehat{M}-M$, and $Z=X-M$. By the basic inequality, we have:
$\forall x\in\mathbb{R}^{n},\qquad x=\Pi x+\Gamma^{\dagger}\Gamma x$
Concentration of (A).
By Cauchy-Schwarz, it is immediate to see that:
$\langle\Pi Z,\Pi\Delta\rangle\leq\|\Pi Z\|_{F}\|\Pi\Delta\|_{F}.$
By Lemma 14, with probability at least $1-o(n^{-1})$:
$(A)\leq 2\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}\|\Delta\|_{F}$
Concentration of (B).
We have:
Then, on $\mathcal{A}$, we have:
Concentration
We thus have:
$\begin{split}\|\Delta\|_{F}^{2}&\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}\|\Delta\|_{F}+\frac{3\rho}{2}\|(\Gamma\Delta)_{S\cdot}\|_{21}\\ &\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}\|\Delta\|_{F}+\frac{3\rho}{2}\sqrt{s}\|\Gamma\Delta\|_{F}\\ &\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}\|\Delta\|_{F}+\frac{3\rho}{2}\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\|\Delta\|_{F}\end{split}$
$\begin{split}\|\Delta\|_{F}&\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}+\frac{3\rho}{2}\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\\ \|\Delta\|_{F}&\leq 4\sqrt{C_{1}Kn_{C}{\frac{\log(n)}{N}}}+6\rho(\Gamma)C_{2}\sqrt{\frac{K\log(n)}{N}}\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\\ &\leq C\left(\sqrt{n_{C}}+\rho(\Gamma)\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\right)\sqrt{\frac{K\log(n)}{N}}\end{split}$
[…]
$\begin{split}\max\{\|\sin\Theta(U,\widehat{U})\|_{F},\|\sin\Theta(V,\widehat{V})\|_{F}\}&\leq\frac{C}{\lambda_{K}(M)}\sqrt{\frac{K\log(n)}{N}}\left(\sqrt{n_{C}}+\rho(\Gamma)\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\right)\\ &\leq CK\sqrt{\frac{\log(n)}{nN}}\left(\sqrt{n_{C}}+\rho(\Gamma)\sqrt{s}\sqrt{\lambda_{\max}(\Gamma)}\right)\end{split}$
This concludes our proof. ∎
We observe first that the error bound for one-step graph-aligned denoising has better rate than the one with no regularization, $O(K\sqrt{\log(n)/N})$, provided in Klopp et al. (2021). We also note that the rate of one-step denoising and GpLSI is equivalent up to a constant. Although the dependency of the error on parameters $n,p,K,$ and $N$ is the same for both methods, our empirical studies in Section 3 reveal that GpLSI still achieves lower errors compared to one-step denoising.
Appendix D Analysis of the Estimation of $W$ and $A$